I'm needing some help with the following problem. Perform the indicated operation and simplify.
(s+t)(s^2-st+t^2)
thanks!

Answer by feliz1965(151)

(s+t)(s^2-st+t^2)
We multiply s by every member of the trinomial(s^2-st+t^2), multiply t by every member of the same trinomial(s^2-st+t^2) and then we finally combine like terms.
s times s^2 = s^3
s times -st = -s^2t
s times t^2 = st^2
===================
We now do the same thing with t.
t times s^2 = s^2t
t times -st = -st^2
t times t^2 = t^3
===================
We now combine everything:

s^3 -s^2t + st^2 + s^2t -st^2 + t^3 =
s^3 + t^3 is the final answer.